OpenStudy (anonymous):

Al's Hatshop bought a hat for $95. Al wants to mark up the hat by 55 percent of the selling price. What should the selling price of the hat be?

OpenStudy (anonymous):

The eventual price: original price + 55% of original price E = 95 + (0.55)(95) = (1.55)(95)